### v1

# The Balitsky-Kovchegov equation for dipole gluon density in the momentum space

Abstract
In this note I review the derivation of the BK equation for the dipole gluon density in momentum space, starting from the BK equation in position space. The note contains no new results; its purpose is to consolidate derivations and formulas scattered across the literature and to establish a consistent notation.
The Balitsky-Kovchegov evolution equation for the dipole amplitude [Balitsky:1995ub, Kovchegov:1999ua] reads
$\frac{\partial}{\partial\ln 1/x}N(\vec{x}_{T},\vec{y}_{T};x)=\frac{\alpha_{s}N_{c}}{2\pi}\int\!d^{2}z_{T}\,\frac{(\vec{x}_{T}-\vec{y}_{T})^{2}}{(\vec{x}_{T}-\vec{z}_{T})^{2}(\vec{y}_{T}-\vec{z}_{T})^{2}}\,\Bigg\{N(\vec{x}_{T},\vec{z}_{T};x)+N(\vec{y}_{T},\vec{z}_{T};x)\\ -N(\vec{x}_{T},\vec{y}_{T};x)-N(\vec{x}_{T},\vec{z}_{T};\eta)N(\vec{z}_{T},\vec{y}_{T};x)\Bigg\}\,.$ (1)
Let us introduces new variables, the dipole size
$\vec{r}_{T}=\frac{\vec{x}_{T}-\vec{y}_{T}}{2}\,$, (2)
and the impact parameter
$\vec{b}_{T}=\frac{\vec{x}_{T}+\vec{y}_{T}}{2}\,.$ (3)
From the position space BK equation one can obtain equation for Transverse Momentum Dependent dipole gluon density. The steps outlined below will closely follow the results obtained in [Kutak:2003bd, Bartels:2007dm].
We will use $k^{2}=|\vec{k}_{T}|^{2}$ as an argument of gluon density. This is the standard notation used in the discussion of angular averaged distributions.
Let’s define the following Fourier transforms [Kimber:2001nm, Kutak:2003bd]
${\cal F}(x,k^{2})=\frac{N_{c}}{\alpha_{s}(2\pi)^{3}}\int d^{2}b\int d^{2}re^{ik\cdot r}\nabla^{2}_{r}\,N(r,b,x)$ (4)
$\Phi(x,k^{2})=\frac{1}{2\pi}\int d^{2}b\int\frac{d^{2}r}{r^{2}}e^{ik\cdot r}N(r,b,x)$ (5)
The function ${\cal F}(x,k^{2})$ is the dipole gluon density. Here the function $\Phi(x,k^{2})$ palyes a role of auxiliary quantity for which the BK equation has a simple form with local nonlinear term as will be shown below. The explicit relation between the two functions ${\cal F}(x,k^{2})$ and $\Phi(x,k^{2})$ is
${\cal F}(x,k^{2})=\frac{N_{c}}{4\alpha_{s}\pi^{2}}k^{2}\nabla^{2}_{k}\Phi(x,k^{2}),\,\,\,\,\Phi(x,k^{2})=\frac{\alpha_{s}\pi^{2}}{N_{c}}\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\ln\frac{l^{2}}{k^{2}}{\cal F}(x,l^{2}).$ (6)
It has been shown in [Kovchegov:1999ua] that once one considers homogenous infinite nucleus the $b$ integral in the eq. (1) can be factorized and enters only via initial condition
$\Phi(x,k^{2})=\frac{1}{2\pi}\int d^{2}b\int\frac{d^{2}r}{r^{2}}e^{ik\cdot r}N(r,b,x)=\frac{1}{2\pi}\int d^{2}b\int\frac{d^{2}r}{r^{2}}e^{ik\cdot r}N(r,x)=\int d^{2}b\,\Phi_{b}(x,k^{2})$ (7)
the function $\Phi_{b}(x,k^{2})$ obeys the following equation
$\begin{split}\frac{\partial\Phi_{b}(x,k^{2})}{\partial\ln 1/x}=\overline{\alpha}_{s}\int_{0}^{\infty}\frac{dl^{2}}{l^{2}}\bigg[\frac{l^{2}\Phi_{b}(x,l^{2})-k^{2}\Phi_{b}(x,k^{2})}{|k^{2}-l^{2}|}+\frac{k^{2}\Phi_{b}(x,k^{2})}{\sqrt{(4l^{4}+k^{4})}}\bigg]-\overline{\alpha}_{s}\Phi^{2}_{b}(x,k^{2}).\end{split}$ (8)
where $\overline{\alpha}=N_{c}\alpha_{s}/\pi$ and which is obtained by Fourier transform of eq. (5) [Kovchegov:1999ua, Marquet:2005zf].
Very explicit presentation of the transform can be found in H. Mantyssari master thesis Balitzky-Kovchegov equation.
The equation for ${\cal F}(x,k^{2})$ is obtained from eq. (8) in several steps. First inserting in the nonlinear part of it relation expressing $\Phi_{b}(x,k^{2})$ in terms of ${\cal F}_{b}(x,k^{2})$ and acting on the whole equation with the operator [Kutak:2012qk]:.
$\frac{N_{c}}{4\,\alpha_{s}\,\pi^{2}}\,k^{2}\,\nabla^{2}_{\!k}\,\Phi_{b}(x,k^{2})\,.$ (9)
that transforms $\Phi_{b}(x,k^{2})$ to ${\cal F}_{b}(x,k^{2})$. For method based on Mellin transform see [Kutak:2006rn].
The linear part of the equation can be written in the form of a differential operator. We have:
$\frac{\partial\Phi_{b}(x,k^{2})}{\partial\ln 1/x}=\alpha_{s}\,\chi\!\left(-\frac{\partial}{\partial\ln k^{2}}\right)\,\Phi_{b}\!\left(x,k^{2}\right)-\bar{\alpha}_{s},\Phi^{2}_{b}\!\left(x,k^{2}\right),$ (10)
where
$\chi(\gamma)=2\psi(1)-\psi(\gamma)-\psi(1-\gamma)$ (11)
is the characteristic function of the BFKL kernel. We represent the BFKL kernel of the BK equation as a power series around $\gamma_{c}=0.373$:
$\chi\!\left(-\partial_{\ln k^{2}}\right)\Phi_{b}(x,k^{2})$ $=\Bigg[\chi(\gamma_{c})+\bigl(-\partial_{\ln k^{2}}-\gamma_{c}\bigr)\chi^{\prime}(\gamma_{c})+\frac{1}{2!}\bigl(-\partial_{\ln k^{2}}-\gamma_{c}\bigr)^{2}\chi^{\prime\prime}(\gamma_{c})+\cdots\Bigg]\Phi_{b}(x,k^{2}),$ (12)
The laplacian can be pulled through the derivative and acts directly on $\Phi_{b}$. After that the full kernel can be restored. The nonlinear term requires explicit evaluation of laplacian acting on the integral operator.
$\frac{\partial\mathcal{F}_{b}(x,k^{2})}{\partial\ln(1/x)}$ $=\frac{N_{c}\alpha_{s}}{\pi}\int_{0}^{\infty}\frac{dl^{2}}{l^{2}}\Bigg[\frac{l^{2}\mathcal{F}_{b}(x,l^{2})-k^{2}\mathcal{F}_{b}(x,k^{2})}{\lvert k^{2}-l^{2}\rvert}+\frac{k^{2}\mathcal{F}_{b}(x,k^{2})}{\sqrt{4l^{4}+k^{4}}}\Bigg]$ (13)
$\quad-\frac{1}{4}\pi\,\alpha_{s}^{2}\,k^{2}\,\nabla^{2}_{\!k}\left[\left(\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\ln\!\frac{l^{2}}{k^{2}}\,\mathcal{F}_{b}(x,l^{2})\right)^{2}\right].$
The BK equation in this form has been for the first time presented in [Bartels:2006ea, Kutak:2006rn]. For another derivation see [Nikolaev:2006za]. Taking the above into account in order to obtain the integrated over impact parameter gluon density one needs to assume some impact parameter dependence since one needs to perform the integration over it to calculate gluon density. For cylinder like target, one may use the following ansatz for factorization [Kutak:2003bd]
${\cal F}_{b}(x,k^{2})={\cal F}(x,k^{2})S(b)$ (14)
with normalization conditions
$\int d^{2}{\bf b}\,S(b)=1,\,\,\,\int d^{2}{\bf b}\,S^{2}(b)=\frac{1}{\pi R^{2}}$ (15)
where $S(b)$ is the profile function $S(b)=\theta(R-b)/\pi R^{2}$.
$\frac{\partial\mathcal{F}(x,k^{2})}{\partial\ln(1/x)}=\frac{N_{c}\alpha_{s}}{\pi}\int_{0}^{\infty}\frac{dl^{2}}{l^{2}}\left[\frac{l^{2}\mathcal{F}(x,l^{2})-k^{2}\mathcal{F}(x,k^{2})}{\lvert k^{2}-l^{2}\rvert}+\frac{k^{2}\mathcal{F}(x,k^{2})}{\sqrt{4l^{4}+k^{4}}}\right]-\frac{\alpha_{s}^{2}k^{2}}{4R^{2}}\,\nabla^{2}_{\!k}\left(\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\ln\!\frac{l^{2}}{k^{2}}\,\mathcal{F}(x,l^{2})\right)^{2}.$ (16)
In the above $\nabla_{k^{2}}^{2}=4\left(\frac{\partial}{\partial k^{2}}+k^{2}\frac{\partial^{2}}{\partial(k^{2})^{2}}\right)$. Let’s calculate the first and second derivative of the nonlinear term in eq. (16)
$\frac{\partial}{\partial k^{2}}\left(\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\ln\!\frac{l^{2}}{k^{2}}\,\mathcal{F}(x,l^{2})\right)^{2}$ $=-\frac{2}{k^{2}}\left(\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\ln\!\frac{l^{2}}{k^{2}}\,\mathcal{F}(x,l^{2})\right)\left(\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\mathcal{F}(x,l^{2})\right).$ (17)
$\frac{\partial^{2}}{\partial(k^{2})^{2}}\left(\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\ln\!\frac{l^{2}}{k^{2}}\,\mathcal{F}(x,l^{2})\right)^{2}$ $=2\Bigg[\left(\frac{1}{k^{2}}\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\mathcal{F}(x,l^{2})\right)^{2}$ (18)
$\qquad\quad+\frac{1}{(k^{2})^{2}}\left(\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\ln\!\frac{l^{2}}{k^{2}}\,\mathcal{F}(x,l^{2})\right)\left(\mathcal{F}(x,k^{2})+\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\mathcal{F}(x,l^{2})\right)\Bigg].$
Combining the second order and first order derivatives together we get
$\nabla^{2}_{\!k}\left[\left(\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\ln\!\frac{l^{2}}{k^{2}}\,\mathcal{F}(x,l^{2})\right)^{2}\right]$ $=\frac{8}{k^{2}}\Bigg\{\left(\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\mathcal{F}(x,l^{2})\right)^{2}$ (19)
$\qquad\quad+\;\mathcal{F}(x,k^{2})\left(\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\ln\!\frac{l^{2}}{k^{2}}\,\mathcal{F}(x,l^{2})\right)\Bigg\}.$
Inserting eq. (19) in the in the eq. (16) we obtain [Kutak:2003bd, Bartels:2007dm]
$\frac{\partial\mathcal{F}(x,k^{2})}{\partial\ln(1/x)}$ $=\frac{N_{c}\alpha_{s}}{\pi}\int_{0}^{\infty}\frac{dl^{2}}{l^{2}}\left[\frac{\,l^{2}\mathcal{F}(x,l^{2})-k^{2}\mathcal{F}(x,k^{2})\,}{\lvert k^{2}-l^{2}\rvert}+\frac{k^{2}\mathcal{F}(x,k^{2})}{\sqrt{4l^{4}+k^{4}}}\right]$
$\quad-\frac{2\,\alpha_{s}^{2}}{R^{2}}\Bigg\{\left[\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\mathcal{F}(x,l^{2})\right]^{2}+\mathcal{F}(x,k^{2})\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\ln\!\left(\frac{l^{2}}{k^{2}}\right)\mathcal{F}(x,l^{2})\Bigg\}.$
For generalizations of the equation to take into account DGLAP terms and kinematical constraint corrections [Kwiecinski:1997ee, Kutak:2003bd, Kutak:2025wjy] it is convenient to rewrite the equation as a double integral equation:
$\mathcal{F}(x,k^{2})={}$ $\;\mathcal{F}_{0}(x,k^{2})+\bar{\alpha}_{s}\int_{x/x_{0}}^{1}\frac{dz}{z}\int_{0}^{\infty}\frac{dl^{2}}{l^{2}}\,\Bigg[\frac{\,l^{2}\mathcal{F}\!\left(\tfrac{x}{z},l^{2}\right)-k^{2}\mathcal{F}\!\left(\tfrac{x}{z},k^{2}\right)\,}{\lvert k^{2}-l^{2}\rvert}+\frac{k^{2}\mathcal{F}\!\left(\tfrac{x}{z},k^{2}\right)}{\sqrt{4l^{4}+k^{4}}}\Bigg]$ (20)
$\;-\frac{2\bar{\alpha}_{s}^{2}\pi^{2}}{N_{c}^{2}R^{2}}\int_{x/x_{0}}^{1}\frac{dz}{z}\,\Bigg\{\Bigg[\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\mathcal{F}\!\left(\tfrac{x}{z},l^{2}\right)\Bigg]^{2}+\mathcal{F}\!\left(\tfrac{x}{z},k^{2}\right)\int_{k^{2}}^{\infty}\frac{dl^{2}}{l^{2}}\,\ln\!\Bigg(\frac{l^{2}}{k^{2}}\Bigg)\,\mathcal{F}\!\left(\tfrac{x}{z},l^{2}\right)\Bigg\}.$
In this form eq. (20) of the BK one can directly see why as $k^{2}$ gets smaller the gluon density gets suppressed for small $k^{2}$ giving rise to soliton like shape. Namely the integral over the nonlinear term is performed over longer domain what suppresses more the linear part of the equation.
The eq. (20) can be also obtained directly in the momentum space [Kutak:2006rn]. The derivation is based on considering propagation of regeizzed gluons in the channel taking into account bootstrap and $2\to 2$ and $2\to 4$ transition kernels projected onto color singlets. One also assumes large number of colors approximation.